OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -3, -5, -7, -9, ...

OpenStudy (anonymous):

I know just add -2 but the answer choices are confusing me an= -3-2 an= -3 +2(n) an= -3 + -2( n+1) an= -3 + -2( n-1)

OpenStudy (welshfella):

nthe term of an arithmetic series is a1 + (n - 1)d where a1 = first term and d = common difference

OpenStudy (anonymous):

Thanks @welshfella

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

its worth remembering that formula